Find z* for each of these situations, taking into account whether the test is one-sided or two-sided.   Then find the p-value, indicating its relation to alpha. Finally, determine if the hypothesis test would lead to rejection of the null.

  1. Test statistic(z) = 2.10,              Ho: p = .10, Ha : p > .10,               α = .05
Answer #1

Solution :

This is the right tailed test .

The null and alternative hypothesis is

H0 : p =2.10

Ha : p > 2.10

Test statistic = z =2.10

P(z > 2.10) = 1 - P(z <2.10 ) = 1-0.9821=0.0179

P-value = 0.0179

= 0.05

P-value <

Reject the null hypothesis .